The probability that a student is admitted to MCBS for bachelor is 0.6. If 10 students from the same school apply, find the probability that (i) at most 5 are admitted. (ii) at least 7 are rejected


1
Expert's answer
2021-08-25T18:26:50-0400

Let "X=" the number of admitted students: "X\sim Bin(n, p)."

Given "p=0.6, n=10, q=1-p=0.4."

(i)


"P(X\leq5)=P(X=0)+P(X=1)+P(X=2)"

"+P(X=3)+P(X=4)+P(X=5)"

"=\dbinom{10}{0}(0.6)^0(0.4)^{10-0}+\dbinom{10}{1}(0.6)^1(0.4)^{10-1}"

"+\dbinom{10}{2}(0.6)^2(0.4)^{10-2}+\dbinom{10}{3}(0.6)^3(0.4)^{10-3}"

"+\dbinom{10}{4}(0.6)^4(0.4)^{10-4}+\dbinom{10}{5}(0.6)^5(0.4)^{10-5}"

"\approx0.3667"


(ii)


"P(X\leq3)=P(X=0)+P(X=1)+P(X=2)"

"+P(X=3)=\dbinom{10}{0}(0.6)^0(0.4)^{10-0}"

"+\dbinom{10}{1}(0.6)^1(0.4)^{10-1}+\dbinom{10}{2}(0.6)^2(0.4)^{10-2}"

"+\dbinom{10}{3}(0.6)^3(0.4)^{10-3}\approx0.0548"